# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v1

@@ III Improved Lower Bound on Access Cost
$|B\cup N|\leqslant\left\lfloor\frac{\Delta_{1}}{r+\delta-1}\right\rfloor\cdot(r+\delta-1)\leqslant\Delta_{1},$
where the first inequality holds by the conditions (i) and (ii). Note that
$|(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\setminus(B\cup N)|\geqslant|\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}|-|B\cup N|\geqslant|\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}|-\Delta_{1}=d-1.$
@@
Therefore, we can obtain an $(n_{F}-d+1)$-subset $S\subseteq\mathcal{F}$ by removing any $d-1$ elements in $(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\setminus(B\cup N)$ from $\mathcal{F}$. Note that both $B\cup N$ and $\mathcal{F}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$ are subsets of $S$. Thus, by $n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$, we have $|\tilde{\mathcal{C}}|_{S}|=q^{k}$.
@@
Let $S_{1}=S\cap(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$. Then, $|S_{1}|=|(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})|-d+1=\Delta_{1}$. Since $B\subseteq(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$ and $B\cup N\subseteq S$, we have $B\subseteq S_{1}$. Moreover, by $N\cap B=\emptyset$ from the condition (iii), we have $N\subseteq S\setminus B$. Thus, $\bm{v}|_{B}$ can be recovered from the symbols $\bm{v}_{N}$ and then from the symbols in $\bm{v}_{S\setminus B}$.
@@
Recall that $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}=\bm{v}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$. Thus, by $(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\cap S=S_{1}$, the symbols in $\bm{v}|_{S}$ that are inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$ originate from $\bm{c}_{1}|_{S_{1}}$. This implies that the non-constant symbols in $\bm{v}|_{S\setminus B}$ are determined by $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$ and $\bm{c}_{1}|_{S_{1}\setminus B}$. Therefore, $|\tilde{\mathcal{C}}|_{S}|$ is upper bounded by
$q^{|\mathcal{R}_{1,1}\cup(S_{1}\setminus B)|}=q^{|\mathcal{R}_{1,1}|+|S_{1}|-|B|},$
@@
where the equality follows from $S_{1}\setminus B\subseteq\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}$, $(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\cap\mathcal{R}_{1,1}=\emptyset$, and $B\subseteq S_{1}$. This leads to $|\mathcal{R}_{1,1}|+|S_{1}|-|B|\geqslant k$, which further implies that $|\mathcal{R}_{1,1}|\geqslant k-\Delta_{1}+|B|$.
@@
For the second part, we consider the case such that $d>n_{I}-k+1$. In this case, we claim that one must have $\Delta_{1}\leqslant 0$. Otherwise, assume that $\Delta_{1}>0$ and let $D\subseteq\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}$ such that $|D|=d-1$. Let $S=\mathcal{F}\setminus D$. By $|S|=n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$, we have $|\tilde{\mathcal{C}}|_{S}|=q^{k}$. Note that for each $\bm{v}\in\tilde{\mathcal{C}}$, the non-constant symbols of $\bm{v}|_{S}$ are either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus D}$ or determined by $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$. Therefore, we have $|\tilde{\mathcal{C}}|_{S}|=q^{k}\leqslant q^{|(\mathcal{U}_{1,1}\setminus D)\cup\mathcal{R}_{1,1}|}$. Since $D\subseteq\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}$, we have $(\mathcal{U}_{1,1}\setminus D)\cup\mathcal{R}_{1,1}=(\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1})\setminus D$ and $|(\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1})\setminus D|=|\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}|-|D|$. This leads to
$|(\mathcal{U}_{1,1}\setminus D)\cup\mathcal{R}_{1,1}|=|\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}|-|D|\geqslant k.$
On the other hand, since $\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}\subseteq\{1\}\times[n_{I}]$, we have
$k\leqslant|\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}|-|D|\leqslant n_{I}-d+1.$
@@
This implies that $d\leqslant n_{I}-k+1$, which is a contradiction. Thus, when $d>n_{I}-k+1$, we always have $\Delta_{1}\leqslant 0$ and therefore, $|\mathcal{R}_{1,1}|\geqslant k$, which completes the proof. ∎
Remark 2:
For $\delta=2$, the condition (ii) can be satisfied only under the case $|B_{i}|=1$ for $1\leqslant i\leqslant\tau$.
@@
The lower bound in Lemma 5 is affected by the size of the subset $|B|$. In the following, we show that there always exists a set $B$ with large cardinality.
@@
Assume that $\mathcal{C}_{F}$ is a linear $[n_{F},tk]_{q}$ LRC with $(r,\delta)_{a}$-locality. For any $\tilde{B}\subseteq[n_{F}]$, there is a subset $B\subseteq\tilde{B}$ of size at least $(\delta-1)\lfloor\frac{|\tilde{B}|}{r+\delta-1}\rfloor$ such that
@@
    $B=\cup_{1\leqslant t\leqslant\tau}B_{t}$ with $B_{t_{1}}\cap B_{t_{2}}=\emptyset$, $t_{1}\neq t_{2}$ and $\tau\leqslant\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$;
@@
    For each $1\leqslant t\leqslant\tau$ and $j_{1},j_{2}\in B_{t}$ with $j_{1}\neq j_{2}$, we may set $\mathrm{Loc}(j_{1})$ and $\mathrm{Loc}(j_{2})$ as the same set, i.e., $\mathrm{Loc}(j_{1})=\mathrm{Loc}(j_{2});$
  • (iii)
    For each $1\leqslant t\leqslant\tau$ and $j\in B_{t}$, we have one possible $\mathrm{Loc}(j)$ satisfying
    $\mathrm{Loc}(j)\cap B=\emptyset.$
@@
We may find a set $B$ by the following process. For $\tau=1$ and $i_{1}\in\tilde{B}$, let $\tilde{B}_{i_{1}}=\tilde{B}\cap S_{i_{1}}$,
$B_{i_{1}}=\begin{cases}B^{*}\subseteq\tilde{B}_{i_{1}},&\text{ if }|\tilde{B}_{i_{1}}|\geqslant\delta-1\text{ and }|B^{*}|=\delta-1,\\ \tilde{B}_{i_{1}},&\text{ if }|\tilde{B}_{i_{1}}|<\delta-1.\end{cases}$ (7)
@@
$\tilde{B}\leftarrow\tilde{B}\setminus S_{i_{1}}.$ (8)
For $1<\tau\leqslant|\tilde{B}|$, if $|\tilde{B}|\neq\emptyset$ then we set $\tilde{B}_{i_{\tau}}=\tilde{B}\cap S_{i_{\tau}}$ for a $i_{\tau}\in\tilde{B}$
$B_{i_{\tau}}=\begin{cases}B^{*}\subseteq\tilde{B}_{i_{\tau}},&\text{ if }|\tilde{B}_{i_{\tau}}|\geqslant\delta-1\text{ and }|B^{*}|=\delta-1,\\ \tilde{B}_{i_{\tau}},&\text{ if }|\tilde{B}_{i_{\tau}}|<\delta-1,\end{cases}$ (9)
and
$\tilde{B}\leftarrow\tilde{B}\setminus S_{i_{\tau}}.$ (10)
For $1<\tau\leqslant|\tilde{B}|$, if $\tilde{B}=\emptyset$, then set
$B=\bigcup_{1\leqslant j\leqslant\tau-1}B_{i_{j}}.$ (11)
By (8) and (10), we have $B$ satisfies the condition (i). By (7) and (9), we have $B$ fulfills the condition (ii) since we can find a $\textrm{Loc}(\ell)\subseteq S_{i_{j}}\setminus B_{i_{j}}$ such that the code symbol $c_{\ell^{\prime}}$ can be recovered by the code symbols with index $\textrm{Loc}(\ell)$ for all $\ell^{\prime}\in B_{i_{j}}$. For the condition (iii), we have
$\begin{split}|B|=&\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|<\delta-1}|B_{i_{j}}|+\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|\geqslant\delta-1}|B_{i_{j}}|\\ =&T+\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|\geqslant\delta-1}(\delta-1)\\ \geqslant&T+(\delta-1)\left\lfloor\frac{|\tilde{B}|-T}{r+\delta-1}\right\rfloor\\ \geqslant&(\delta-1)\left\lfloor\frac{|\tilde{B}|}{r+\delta-1}\right\rfloor,\end{split}$ (12)
where
$T=\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|<\delta-1}|B_{i_{j}}|$
and the first inequality holds by the fact that in each round the set $\tilde{B}$ reduces $|B_{i_{j}}|$ for the case $|\tilde{B}_{i_{j}}|<\delta-1$ and at most $|S_{i_{j}}|\leqslant r+\delta-1$ when $|\tilde{B}_{i_{j}}|\geqslant\delta-1$. This completes the proof. ∎
@@
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$ where $\delta\geqslant 2$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$. Then, the write access cost is
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0\\ t\left(k-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right),&\mbox{otherwise}.\end{cases}$
@@
$|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|\leqslant|\mathcal{U}_{i,1}|\leqslant n_{F}-d-(t-1)k+1-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1),$

### v2

@@ III Improved Lower Bound on Access Cost
$|B\cup N|\leqslant|B|+|\mathcal{A}_{1}|r\leqslant\left\lfloor\frac{\Delta_{1}}{r+\delta-1}\right\rfloor\cdot(r+\delta-1)\leqslant\Delta_{1},$
where the first inequality holds by $|S_{t}|\leqslant r+\delta-1$ for $t\in[\tau]$, and the second inequality comes from condition (i) that $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for $1\leqslant t_{2}<t_{1}\leqslant\tau$ which implies $|\mathcal{A}_{1}|\leqslant\lfloor\frac{\Delta_{1}}{r+\delta-1}\rfloor$ by $|B|\leqslant(\delta-1)\lfloor\frac{\Delta_{1}}{r+\delta-1}\rfloor$. Note that
$|(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})\setminus(B\cup N)|\geqslant|\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}|-|B\cup N|\geqslant|\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}|-\Delta_{1}=d-1.$
@@
Therefore, we can obtain an $(n_{F}-d+1)$-subset $M\subseteq\mathcal{F}_{1}$ by removing any d1d-1 elements in $(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})\setminus(B\cup N)$ from $\mathcal{F}_{1}$. Note that both $B\cup N$ and $\mathcal{F}_{1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})$ are subsets of $M$. Thus, by $n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$, we have $|\tilde{\mathcal{C}}|_{M}|=q^{k}$.
Claim 1: We claim that the symbols $\bm{v}|_{B_{1}\cup\cdots\cup B_{\ell}}$ can be recovered by the symbols $\bm{v}|_{M\setminus(B_{1}\cup\cdots\cup B_{\ell})}$ for $1\leqslant\ell\leqslant\tau$. The claim can be proved by induction as follows. Denote $\overline{\mathcal{U}}\triangleq\mathcal{U}_{2,1}\cup\cdots\cup\mathcal{U}_{t,1}$. If $\ell=1\in\mathcal{A}_{2}$, then $|S_{1}\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})|<\delta-1$ by (13). Since (9) and $S_{1}$ is a repair set in $\mathcal{C}_{F}$, then $S_{1}\subseteq\mathcal{F}_{1}=\mathcal{U}_{1,1}\cup\overline{\mathcal{U}}\cup\mathcal{W}_{1}$. Furthermore, we have
$S_{1}\cap((\mathcal{U}_{1,1}\cap\mathcal{D}_{1,1})\cup\overline{\mathcal{U}}\cup\mathcal{W}_{1})$
$=$~ $S_{1}\cap(\mathcal{F}_{1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}))$
and
$|S_{1}\cap((\mathcal{U}_{1,1}\cap\mathcal{D}_{1,1})\cup\overline{\mathcal{U}}\cup\mathcal{W}_{1})|$
$=$~ $|S_{1}\cap(\mathcal{F}_{1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}))|$
$=$~ $|S_{1}|-|S_{1}\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})|$
$>$~ $|S_{1}|-\delta+1.$
@@
Thus, by $(r,\delta)$-locality and the fact $S_{1}$ is a repair set, we obtain that $\bm{v}|_{B_{1}}$ can be recovered by $\bm{v}|_{\mathcal{F}_{1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})}$ and then can be recovered by $\bm{v}|_{M\setminus B_{1}}$ since F_1 ∖(U_1,1 ∖D_1,1)⊆M, B_1⊆M, B_1⊆U_1,1 ∖D_1,1, i.e., B_1∩(F_1 ∖(U_1,1 ∖D_1,1))=∅. If $\ell=1\in\mathcal{A}_{1}$, the repair set $S_{1}\subseteq B\cup N\subseteq M$, therefore, the symbols $\bm{v}|_{B_{1}}$ can be recovered by the symbols $\bm{v}|_{M\setminus B_{1}}$ by condition (ii) that $|B_{1}|\leqslant\delta-1$. In both subcases, we have $v|_{S_{1}}$ can also be recovered by $\bm{v}|_{M\setminus B_{1}}$.
Assume that the symbols $\bm{v}|_{B_{1}\cup\dots\cup B_{\ell}}$ and $\bm{v}|_{S_{1}\cup\dots\cup S_{\ell}}$ can be recovered by the symbols $\bm{v}|_{M\setminus(B_{1}\cup\dots\cup B_{\ell})}$ for $1\leqslant\ell<\tau$. Consider the case $\ell+1$. Since condition (i) that $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$, then the symbols $\bm{v}|_{B_{1}\cup\dots\cup B_{\ell}}$ and $\bm{v}|_{S_{1}\cup\dots\cup S_{\ell}}$ can be recovered by the symbols $\bm{v}|_{M\setminus(B_{1}\cup\dots\cup B_{\ell+1})}$ by our assumption. It is sufficient to show that the symbols $\bm{v}|_{B_{\ell+1}}$ and $\bm{v}|_{S_{\ell+1}}$ can be recovered by the symbols $\bm{v}|_{M\setminus B_{\ell+1}}$ since the symbols $\bm{v}|_{B_{1}\cup\dots\cup B_{\ell}}$ are recoverable in this case.
For the subcase $\ell+1\in\mathcal{A}_{2}$, by (13), we have
$\left|\left(S_{\ell+1}\setminus\bigcup_{i\in[\ell]}S_{i}\right)\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})\right|<\delta-1,$ (14)
which implies
$\left|\left(S_{\ell+1}\cap\left(M\setminus B_{\ell+1}\right)\right)\cup\left(\bigcup_{i\in[\ell]}(S_{i}\cap S_{\ell+1})\right)\right|$
$\geqslant$~ $\left|(S_{\ell+1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}))\cap(M\setminus B_{\ell+1}))\cup\left(\bigcup_{i\in[\ell]}(S_{i}\cap S_{\ell+1})\right)\right|$
$=$~ $\left|(S_{\ell+1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}))\cup\left(\bigcup_{i\in[\ell]}(S_{i}\cap S_{\ell+1})\right)\right|$
$=$~ $\left|S_{\ell+1}\setminus\left(\left(S_{\ell+1}\setminus\bigcup_{i\in[\ell]}S_{i}\right)\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})\right)\right|$
$\geqslant$~ $|S_{\ell+1}|-\delta+1,$
where the first equation holds by $S_{\ell+1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})\subseteq M$ and $B_{\ell+1}\subseteq\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}$, i.e., $(S_{\ell+1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}))\subseteq(M\setminus B_{\ell+1})$, the last inequality comes from (14). Then, the symbols $\bm{v}|_{B_{\ell+1}}$ can be recovered by the symbols $\bm{v}|_{M\setminus B_{\ell+1}}$ according to the $(r,\delta)$-locality.
If $\ell+1\in\mathcal{A}_{1}$, then $S_{\ell+1}\subseteq B\cup N\subseteq M$. Since the symbols $\bm{v}|_{S_{1}\cup\dots\cup S_{\ell}}$ are known, we have that
$\left|(S_{\ell+1}\cap(M\setminus B_{\ell+1}))\cup\left(\bigcup_{i\in[\ell]}(S_{i}\cap S_{\ell+1})\right)\right|$
$\geqslant$~ $\left|(S_{\ell+1}\setminus B_{\ell+1})\cup\left(\bigcup_{i\in[\ell]}(S_{i}\cap S_{\ell+1})\right)\right|$
$=$~ $\left|S_{\ell+1}\setminus B_{\ell+1}\right|$
$\geqslant$~ $|S_{\ell+1}|-\delta+1,$
where the first inequality holds by $B_{\ell+1}\subseteq S_{\ell+1}\subseteq B\cup N\subseteq M$, the first equation comes from the condition (i) that $B_{\ell+1}\cap S_{i}=\emptyset$ for $i\in[\ell]$, and the second inequality holds by condition (ii) that $|B_{\ell+1}|\leqslant\delta-1$. Then, the symbols $\bm{v}|_{B_{\ell+1}}$ and $\bm{v}|_{S_{\ell+1}}$ can be recovered by the symbols $\bm{v}|_{M\setminus B_{\ell+1}}$. Therefore, the symbols $\bm{v}|_{B_{1}\cup\cdots\cup B_{\tau}}=\bm{v}|_{B}$ can be recovered by $\bm{v}|_{M\setminus(B_{1}\cup\cdots\cup B_{\tau})}=\bm{v}|_{M\setminus B}$ according to induction.
@@
Let $M_{1}=M\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})$. Then, $|M_{1}|=|(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})|-d+1=\Delta_{1}$. Recall that $B\subseteq\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}$ and $B\cup N\subseteq M$. Since $\bm{v}|_{M\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})}$ are determined by $\bm{c}_{1}|_{\mathcal{D}_{1,1}}$, thus the non-constant symbols in $\bm{v}|_{M\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})}$ and $\bm{v}|_{M_{1}\setminus B}$ (i.e., $\bm{v}|_{M\setminus B}$) are determined by $\bm{c}_{1}|_{\mathcal{D}_{1,1}}$ and $\bm{c}_{1}|_{M_{1}\setminus B}$. Now, by the above Claim 1, we have the non-constant symbols in $\bm{v}|_{M}$ are determined by $\bm{c}_{1}|_{\mathcal{D}_{1,1}}$ and $\bm{c}_{1}|_{M_{1}\setminus B}$. Therefore, $|\tilde{\mathcal{C}}|_{M}|$ is upper bounded by
$q^{|\mathcal{D}_{1,1}\cup(M_{1}\setminus B)|}=q^{|\mathcal{D}_{1,1}|+|M_{1}|-|B|},$
@@
where the equality follows from $M_{1}\setminus B\subseteq\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}$, $(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})\cap\mathcal{D}_{1,1}=\emptyset$, and $B\subseteq M_{1}$. This leads to $|\mathcal{D}_{1,1}|+|M_{1}|-|B|\geqslant k$, which further implies that $|\mathcal{D}_{1,1}|\geqslant k-\Delta_{1}+|B|$.
@@
Case 3: $d>n_{I}-k+1$. We claim that one must have $\Delta_{1}\leqslant 0$. Otherwise, assume that $\Delta_{1}>0$ and let $Q\subseteq\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}$ such that |Q|=d1|Q|=d-1. Let $M=\mathcal{F}_{1}\setminus Q$. By $|M|=n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$, we have $|\tilde{\mathcal{C}}|_{M}|=q^{k}$. Note that for each $\bm{v}\in\tilde{\mathcal{C}}$, the non-constant symbols of $\bm{v}|_{M}$ are either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus Q}$ or determined by $\bm{c}_{1}|_{\mathcal{D}_{1,1}}$. Therefore, we have $|\tilde{\mathcal{C}}|_{M}|=q^{k}\leqslant q^{|(\mathcal{U}_{1,1}\setminus Q)\cup\mathcal{D}_{1,1}|}$. Since $Q\subseteq\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}$, we have $(\mathcal{U}_{1,1}\setminus Q)\cup\mathcal{D}_{1,1}=(\mathcal{U}_{1,1}\cup\mathcal{D}_{1,1})\setminus Q$ and $|(\mathcal{U}_{1,1}\cup\mathcal{D}_{1,1})\setminus Q|=|\mathcal{U}_{1,1}\cup\mathcal{D}_{1,1}|-|Q|$. This leads to
$|(\mathcal{U}_{1,1}\setminus Q)\cup\mathcal{D}_{1,1}|=|\mathcal{U}_{1,1}\cup\mathcal{D}_{1,1}|-|Q|\geqslant k.$
On the other hand, since $\mathcal{U}_{1,1}\cup\mathcal{D}_{1,1}\subseteq\{1\}\times[n_{I}]$, we have
$k\leqslant|\mathcal{U}_{1,1}\cup\mathcal{D}_{1,1}|-|Q|\leqslant n_{I}-d+1.$
@@
This implies that $d\leqslant n_{I}-k+1$, which is a contradiction. Thus, when $d>n_{I}-k+1$, we always have $\Delta_{1}\leqslant 0$ and therefore, $|\mathcal{D}_{1,1}|\geqslant k$, which completes the proof. ∎
@@
The lower bound in Lemma 6 depends critically on the size of the subset $|B|$. Below, we demonstrate that there always exists a set $B$ with large cardinality.
@@
Assume that $\mathcal{C}$ is a linear $n_{F},tk]_{q}$ LRC with $(r,\delta)_{a}$-locality. For any $\tilde{V}\subseteq[n_{F}]$, there is a subset $B=\cup_{1\leqslant t\leqslant\tau}B_{t}\subseteq\tilde{V}$ of size $(\delta-1)\lfloor\frac{|\tilde{V}|}{r+\delta-1}\rfloor$ for a positive integer $\tau$, such that
@@
    For each $1\leqslant t\leqslant\tau$, the symbols $c_{i}$ for all $i\in B_{t}$ are in the same repair set $S_{t}$ of the code $\mathcal{C}$, and $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$;
@@
    For each $1\leqslant t\leqslant\tau$, $|B_{t}|\leqslant\delta-1$, and
    $\left|\left(S_{t}\setminus\bigcup_{1\leqslant i\leqslant t-1}S_{i}\right)\cap\tilde{V}\right|<\delta-1,$ (15)
    if $|B_{t}|<\delta-1$.
@@
We find a subset $V$ of $\tilde{V}$ by the following process. For $\ell=1$ and $i_{1}\in V^{(1)}=\tilde{V}$, let $\tilde{V}_{i_{1}}=V^{(1)}\cap S_{i_{1}}$,
[…]
@@
$V^{(2)}=V^{(1)}\setminus S_{i_{1}}.$ (17)
[…]
and then set
$V^{(\ell+1)}=V^{(\ell)}\setminus S_{i_{\ell}}.$ (19)
[…]
$V=\bigcup_{1\leqslant j\leqslant\ell-1}V_{i_{j}}.$
Thus,
$\begin{split}|V|=&\sum_{1\leqslant j\leqslant\ell-1\atop|\tilde{V}_{i_{j}}|<\delta-1}|V_{i_{j}}|+\sum_{1\leqslant j\leqslant\ell-1\atop|\tilde{V}_{i_{j}}|\geqslant\delta-1}|V_{i_{j}}|\\ =&T+\sum_{1\leqslant j\leqslant\ell-1\atop|\tilde{V}_{i_{j}}|\geqslant\delta-1}(\delta-1)\\ \geqslant&T+(\delta-1)\left\lfloor\frac{|\tilde{V}|-T}{r+\delta-1}\right\rfloor\\ \geqslant&(\delta-1)\left\lfloor\frac{|\tilde{V}|}{r+\delta-1}\right\rfloor,\end{split}$ (20)
where T=∑_1⩽j⩽ℓ-1— ~Vij—¡δ-1 —V_i_j— and the first inequality holds by the fact that in each round the set $\tilde{V}$ reduces $|V_{i_{j}}|$ for the case $|\tilde{V}_{i_{j}}|<\delta-1$ and at most $|S_{i_{j}}|\leqslant r+\delta-1$ when $|\tilde{V}_{i_{j}}|\geqslant\delta-1$.
Denote $\mathcal{V}_{1}\triangleq\{j:|V_{i_{j}}|=\delta-1,j\in[\ell-1]\}$ and $\mathcal{V}_{2}\triangleq\{j:|V_{i_{j}}|<\delta-1,j\in[\ell-1]\}$. On the one hand, if $|\mathcal{V}_{1}|\geqslant\lfloor\frac{|\tilde{V}|}{r+\delta-1}\rfloor$, then set $B$ as the union of any $\lfloor\frac{|\tilde{V}|}{r+\delta-1}\rfloor$ sets in $\{V_{i_{j}}\}_{j\in\mathcal{V}_{1}}$. On the other hand, if $|\mathcal{V}_{1}|<\lfloor\frac{|\tilde{V}|}{r+\delta-1}\rfloor$, by (20), there exists $\hat{V}_{i_{j}}\subseteq V_{i_{j}}$ for $j\in\mathcal{V}_{2}$ such that $B=(\cup_{j\in\mathcal{V}_{1}}V_{i_{j}})\cup(\cup_{j\in\mathcal{V}_{2}}\hat{V}_{i_{j}})$ and $|B|=(\delta-1)\lfloor\frac{|\tilde{V}|}{r+\delta-1}\rfloor$.
Let $B_{j}=V_{i_{j}}$ if $j\in\mathcal{V}_{1}$ and $B_{j}=\hat{V}_{i_{j}}$ if $j\in\mathcal{V}_{2}$. Therefore, $B=\cup_{1\leqslant t\leqslant\tau}B_{t}$ and $\tau\leqslant|\mathcal{V}_{1}|+|\mathcal{V}_{2}|$ since $\hat{V}_{i_{j}}$ could be an empty set for $j\in\mathcal{V}_{2}$. The expression $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$ in the condition (i) comes from (17) and (19). The inequality $|B_{t}|\leqslant\delta-1$ in the condition (ii) holds by (16) and (18). The inequality (15) holds by (16)-(19). Thus, we have $B_{j}$s for $j\in[\tau]$ are the desirable subsets of $B$, which completes the proof. ∎
@@
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$ where $\delta\geqslant 2$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$ and the optimal Hamming distance $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\phi(k,r,\delta)$. Then, the write access cost is
$\rho_{w}\geqslant t\left(d+(t-1)k-1+\phi((t-1)k,r,\delta)\right)-(t-1)n_{F}.$ (21)
Denote $\Delta=n_{F}-2d-(t-1)k+2-\phi((t-1)k,r,\delta)$. Then, the read access cost is
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ t\left(k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right),&\mbox{otherwise}.\end{cases}$
@@
$|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|\leqslant|\mathcal{U}_{i,1}|\leqslant{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n_{F}-d-(t-1)k+1-\phi((t-1)k,r,\delta),}$